OpenStudy (anonymous):

I'm having trouble trying to answer this question. "How can you prove that b - a is the negative of a - b?"

OpenStudy (john_es):

May be this way, \[0=(a-a)+(b-b)=a-a+b-b=(a-b)+(b-a)\Rightarrow b-a=-(a-b)\]